can you help me find the answer to 3( y − 2) − 5 = 2y +1;
Answer by unlockmath(1688)

Hello,
Here we go.
3( y − 2) − 5 = 2y +1 This can be rewritten as:
3y − 6 − 5 = 2y +1 Combine like terms:
3y−11 = 2y +1 Now Add 11 to both sides:
3y=2y+12 Then subtract 2y from both sides:
y=12
Plug in 12 to the original to see if it it correct.
RJ
Check out a new book I wrote at:
www.math-unlock.com